# Multi-Derivative Runge-Kutta Flux Reconstruction for hyperbolic conservation laws

### v1

@@ 6.2 Higher order blending
Since the evolution in first stage is from $t^{n}$ to $t^{n+\frac{1}{2}}$ and in the second from $t^{n}$ to $t^{n+1}$, the procedure remains the smae for both stages. Thus, the details of the construction of an admissibility preserving MUSCL-Hancock flux on the non-uniform non-cell centred subcells involved in the blending scheme can be found in [4].
@@ 7.2.1 Blast
in the domain $[0,1]$. This test was originally introduced in [38] to check the capability of the numerical scheme to accurately capture the shock-shock interaction scenario. The boundaries are set as solid walls by imposing the reflecting boundary conditions at $x=0$ and $x=1$. The solution consists of reflection of shocks and expansion waves off the boundary wall and several wave interactions inside the domain. The numerical solutions give negative pressure if the positivity correction is not applied. With a grid of 400 cells using polynomial degree $N=3$, we run the simulation till the time $t=0.038$ where a high density peak profile is produced. As tested in [4], we compare first order (FO) and MUSCL-Hancock (MH) blending schemes, and TVB limiter with parameter $M=300$ [21] (TVB-300). We compare the performance of limiters in Figure (7) where the approximated density and pressure profiles are compared with a reference solution computed using a very fine mesh. Looking at the peak amplitude and contact discontinuity, it is clear that MUSCL-Hancock blending scheme gives the best resolution, especially when compared with the TVB limiter.
@@ 7.3.1 Double Mach reflection
and take the initial condition to be $\bm{u}_{0}(x,y)=\bm{u}_{b}(x,y,0)$. With $\bm{u}_{b}$, we impose inflow boundary conditions at the left side $\{0\}\times[0,1]$, outflow boundary conditions both on $[0,1/6]\times\{0\}$ and $\{4\}\times[0,1]$, reflecting boundary conditions on $[1/6,4]\times\{0\}$ and inflow boundary conditions on the upper side $[0,4]\times\{1\}$.
@@ 7.3.3 Two Dimensional Riemann problem
(a) (b)
(c) (d)
Figure 14: 2-D Riemann problem, density plots of numerical solution at $t=0.25$ for degree $N=3$ on a $256\times 256$ mesh (a) Trixi.jl, (b) MDRK, (c) Trixi.jl zoomed, (d) MDRK zoomed
@@ 7.3.4 Rayleigh-Taylor instability
The implementation of MDRK with source terms is based on [2] where an approximate Lax-Wendroff procedure is also applied on the source term. The following description is taken from [20]. The Rayleigh-Taylor instability happens on the interface between fluids with different densities when an acceleration is directed from the heavy fluid to the light one. The instability with fingering nature generates bubbles of light fluid rising into the ambient heavy fluid and spikes of heavy fluid falling into the light fluid. The initial condition of this problem [28] is given as follows
@@
where $a=\sqrt{\gamma p/\rho}$ is the sound speed and $\gamma=5/3$. The computational domain is $[0,0.25]\times[0,1]$. The reflecting boundary conditions are imposed for the left and right boundaries. At the top boundary, the flow variables are set as $(\rho,u,v,p)=(1,0,0,2.5)$. At the bottom boundary, they are $(\rho,u,v,p)=(2,0,0,1)$. The uniform mesh with $64\times 256$ elements is used in the simulation. The density distributions at $t=1.5,1.75,2,2.25,2.5$ are presented in Figure 15. It is a test to check the suitability of higher-order schemes for the capturing of interface instabilities.
@@
Figure 15: Rayleigh-Taylor instability on a $64\times 256$ mesh
@@ 8 Summary and conclusions
This work introduces fourth order multiderivative Runge-Kutta (MDRK) scheme of [18] in the conservative, quadrature free Flux Reconstruction framework to solve hyperbolic conservation laws. The idea is to cast each MDRK stage as an evolution involving a time average flux which is approximated by the Jacobian free Approximate Lax-Wendroff procedure. The numerical flux is carefully computed with accuracy and stability in mind. In particular, the D2 dissipation and EA flux of [3] have been introduced which enhance stable CFL numbers and accuracy for nonlinear problems respectively. The stable CFL numbers are computed using Fourier stability analysis for two commonly used correction functions $g_{\operatorname{Radau}}$ and $g_{2}$, showing the improved CFL numbers. Convergence analysis for non-linear problems was performed which revealed that optimal convergence rates were only shown when using the EA flux. The shock capturing blending scheme of [4] has also been introduced for the MDRK scheme applied at each stage. The scheme is provably admissibility preserving and good at capturing small scale structures. The claims are validated by numerical experiments.
Appendix A Formal accuracy analysis
@@
Further note that, we use the Approximate Lax-Wendroff (Section 4.3) to approximate $\partial_{t}\bm{L}\left(\bm{u}^{n}\right),\partial_{t}\bm{L}\left(\bm{u}^{\ast}\right)$ to $O(\Delta t^{3})$ accuracy and thus we perform an error analysis of an evolution performed as
@@
Starting from $\bm{u}=\bm{u}^{n}$, the exact solution satisfies
@@
Now we will substitute these four equations into (39) and use (40) to obtain the update equation in terms of temporal derivatives on $\bm{u}$. Then, we compare with the Taylor’s expansion of $\bm{u}$ (41) to get conditions for the respective orders of accuracy
@@
We then see that equations (44), (45) become identical, and equations (46), (47) become identical. Simplifying the above equations, we get five equations for the five unknown coefficient.
@@
These coefficients do give the scheme (6).

### v2

@@ 6.2 Higher order blending
Since the evolution in first stage is from $t^{n}$ to $t^{n+\frac{1}{2}}$ and in the second from $t^{n}$ to $t^{n+1}$, the procedure remains the same for both stages. Thus, the details of the construction of an admissibility preserving MUSCL-Hancock flux on the non-uniform non-cell centred subcells involved in the blending scheme can be found in [4].
@@ 7.2.1 Blast wave
in the domain $[0,1]$. This test was originally introduced in [38] to check the capability of the numerical scheme to accurately capture the shock-shock interaction scenario. The boundaries are set as solid walls by imposing the reflecting boundary conditions at $x=0$ and $x=1$. The solution consists of reflection of shocks and expansion waves off the boundary wall and several wave interactions inside the domain. The numerical solutions give negative pressure if the positivity correction is not applied. With a grid of 400 cells using polynomial degree $N=3$, we run the simulation till the time $t=0.038$ where a high-density peak profile is produced. As tested in [4], we compare first order (FO) and MUSCL-Hancock (MH) blending schemes, and TVB limiter with parameter $M=300$ [21] (TVB-300). We compare the performance of limiters in Figure 7 where the approximated density and pressure profiles are compared with a reference solution computed using a very fine mesh. Looking at the peak amplitude and contact discontinuity, it is clear that the MUSCL-Hancock blending scheme gives the best resolution, especially when compared with the TVB limiter.
@@ 7.3.1 Double Mach reflection
and take the initial condition to be $\bm{u}_{0}(x,y)=\bm{u}_{b}(x,y,0)$. With $\bm{u}_{b}$, we impose inflow boundary conditions on the left side $\{0\}\times[0,1]$, outflow boundary conditions both on $[0,1/6]\times\{0\}$ and $\{4\}\times[0,1]$, reflecting boundary conditions on $[1/6,4]\times\{0\}$ and inflow boundary conditions on the upper side $[0,4]\times\{1\}$.
@@ 7.3.3 Two Dimensional Riemann problem
(a) Trixi.jl (b) MDRK
Figure 14: 2-D Riemann problem, density plots of numerical solution at $t=0.25$ for degree $N=3$ on a $256\times 256$ mesh.
@@ 7.3.4 Rayleigh-Taylor instability
It is also a test to check the suitability of higher-order schemes for the capturing of interface instabilities. The implementation of MDRK with source terms is explained in Appendix B where an approximate Lax-Wendroff procedure is also applied on the source term. The following description of this test is from [20]. The Rayleigh-Taylor instability happens on the interface between fluids with different densities when an acceleration is directed from the heavy fluid to the light one. The instability with fingering nature generates bubbles of light fluid rising into the ambient heavy fluid and spikes of heavy fluid falling into the light fluid. The initial condition of this problem [28] is given as follows
@@
where $a=\sqrt{\gamma p/\rho}$ is the sound speed, $\gamma=5/3$ and the computational domain is $[0,0.25]\times[0,1]$. The reflecting boundary conditions are imposed for the left and right boundaries. At the top boundary, the flow variables are set as $(\rho,u,v,p)=(1,0,0,2.5)$. At the bottom boundary, they are $(\rho,u,v,p)=(2,0,0,1)$. The uniform mesh with $64\times 256$ elements is used in the simulation. The density distributions at $t=1.5,1.75,2,2.25,2.5$ are presented in Figure 15 which follows qualitatively the physical picture described above.
@@
Figure 15: Density for the Rayleigh-Taylor instability on a $64\times 256$ mesh
@@ 8 Summary and conclusions
This work introduces the fourth order multiderivative Runge-Kutta (MDRK) scheme of [18] in the conservative, quadrature free, Lax-Wendroff Flux Reconstruction framework to solve hyperbolic conservation laws. The idea is to cast each MDRK stage as an evolution involving a time average flux which is approximated by the Jacobian-free approximate Lax-Wendroff procedure. The D2 dissipation and EA flux of [3] have been introduced for numerical flux computation, which enhance stable CFL numbers and accuracy for nonlinear problems, respectively. The stable CFL numbers are computed using Fourier stability analysis for two commonly used correction functions, $g_{\operatorname{Radau}}$ and $g_{2}$, showing the improved CFL numbers. Numerical convergence studies for non-linear problems were performed, which revealed that optimal convergence rates were only seen when using the EA flux. The shock capturing blending scheme of [4] has also been introduced for the MDRK scheme applied at each stage. Following [4], flux limiting is applied to each MDRK stage to obtain admissibility preservation in means. The scaling limiter of [41] is then applied to obtain admissibility preservation at the solution points. Along with being provably admissibility preserving, the scheme is more capable of capturing small scale structures through the use of Gauss-Legendre solution points and application of MUSCL-Hancock scheme on the subcells. The claims are validated by numerical experiments for compressible Euler’s equations with the modern test suite [20] for high order methods.
Appendix A Derivation of 2-stage, 4-th order scheme
@@
where the coefficients $a_{ij},b_{i}$ need to be determined to achieve fourth order accuracy. Further note that, we use the approximate Lax-Wendroff procedure (Section 4.3) to approximate $\partial_{t}\bm{L}\left(\bm{u}^{n}\right),\partial_{t}\bm{L}\left(\bm{u}^{\ast}\right)$ to $O(\Delta t^{3})$ accuracy and thus we perform an error analysis of an evolution performed as
@@
where $\bm{L}_{\bm{u}}=\partial\bm{L}/\partial\bm{u}$, etc. Starting from $\bm{u}=\bm{u}^{n}$, the exact solution satisfies
@@
Now we will substitute these four equations into (43) and use (44) to obtain the update equation in terms of temporal derivatives on $\bm{u}$. Then, we compare with the Taylor’s expansion of $\bm{u}$ (45) to get conditions for the respective orders of accuracy.
@@
We then see that equations (48), (49) become identical, and equations (50), (51) become identical. Simplifying the above equations, we get five equations for the five unknown coefficients
@@
These coefficients do give the scheme (7) for which the two stage method is fourth order accurate.
Remark 5
Equations (43) justify the choice of finite difference approximations made in Section 4.3. The time derivatives $\partial_{t}\bm{L}(\bm{u}^{n}),\partial_{t}\bm{L}(\bm{u}^{\ast})$ need to be approximated to atleast third order of accuracy for the one-step error to be $O(\Delta t^{5})$ and the overall method to be fourth order accurate, which is why the third order accurate finite difference formulae in time are used in Section 4.3.
Appendix B Treatment of source terms
In this section, we extend the multi-derivative Runge-Kutta (MDRK) Flux Reconstruction scheme to conservation laws with source terms
$\bm{u}_{t}+\bm{f}\left(\bm{u}\right)_{x}=\text{{{s}}}$ (60)
where $\text{{{s}}}=\text{{{s}}}(\bm{u},t,\bm{x})$. The idea is to apply the MDRK scheme of [18] to (60) as follows
$\begin{split}\bm{u}^{\ast}&=\bm{u}^{n}-\frac{\Delta t}{2}\partial_{x}\bm{F}+\frac{\Delta t}{2}\mathbf{S}\\ \bm{u}^{n+1}&=\bm{u}^{n}-\Delta t\partial_{x}\bm{F}^{\ast}+\Delta t\mathbf{S}^{*}\end{split}$
where $\bm{F},\bm{F}^{\ast}$ are the time averaged fluxes (1213) and $\mathbf{S},\mathbf{S}^{*}$ are the time averaged source terms satisfying
$\begin{split}\mathbf{S}&:=\text{{{s}}}\left(\bm{u}^{n}\right)+\frac{1}{4}\Delta t\frac{\partial}{\partial t}\text{{{s}}}\left(\bm{u}^{n}\right)\approx\frac{1}{\Delta t/2}\int_{t^{n}}^{t^{n+\frac{1}{2}}}\text{{{s}}}\\ \mathbf{S}^{*}&:=\text{{{s}}}(\bm{u}^{n})+\frac{1}{6}\Delta t\left(\frac{\partial}{\partial t}\text{{{s}}}(\bm{u}^{n})+2\frac{\partial}{\partial t}\text{{{s}}}(\bm{u}^{\ast})\right)\approx\frac{1}{\Delta t}\int_{t^{n}}^{t^{n+1}}\text{{{s}}}\end{split}$
The approximate Lax-Wendroff procedure is used to approximate time average fluxes and source terms. It is an extension of the procedure in Section 4.3 to solve (60) which we now describe.
First stage.
The time averaged flux $\bm{F}$ and time averaged source term $\mathbf{S}$ are approximated as
$\mathsf{\bm{F}}=\mathsf{f}+\frac{1}{4}\mathsf{f}^{(1)},\qquad\mathsf{\mathbf{S}}={\mathsf{s}}+\frac{1}{4}{\mathsf{s}}^{(1)}$
where
$\mathsf{u}^{(1)}$ $=-\frac{\Delta t}{\Delta x_{e}}\mathsf{D}\mathsf{f}+\Delta t\text{{{s}}}$
$\mathsf{f}^{(1)}$ $=\frac{1}{12}\left[-\bm{f}(\mathsf{u}+2\mathsf{u}^{(1)})+8\bm{f}(\mathsf{u}+\mathsf{u}^{(1)})-8\bm{f}(\mathsf{u}-\mathsf{u}^{(1)})+\bm{f}(\mathsf{u}-2\mathsf{u}^{(1)})\right]$
${\mathsf{s}}^{(1)}$ $=\frac{1}{12}\left[-\text{{{s}}}(\mathsf{u}+2\mathsf{u}^{(1)})+8\text{{{s}}}(\mathsf{u}+\mathsf{u}^{(1)})-8\text{{{s}}}(\mathsf{u}-\mathsf{u}^{(1)})+\text{{{s}}}(\mathsf{u}-2\mathsf{u}^{(1)})\right]$
Second stage.
The time averaged flux $\bm{F}^{\ast}$ and the time averaged source term $\mathbf{S}^{*}$ are approximated as
$\mathsf{\bm{F}^{\ast}}=\mathsf{f}+\frac{1}{6}(\mathsf{f}^{(1)}+2\mathsf{f}^{*(1)}),\qquad\mathsf{\mathbf{S}^{*}}={\mathsf{s}}+\frac{1}{6}({\mathsf{s}}^{(1)}+2{\mathsf{s}}^{*(1)})$
where
$\mathsf{u}^{*(1)}$ $=-\frac{\Delta t}{\Delta x_{e}}\mathsf{D}\mathsf{f}^{*(1)}+{\mathsf{s}}^{*(1)}$
$\mathsf{f}^{*(1)}$ $=\frac{1}{12}\left[-\bm{f}(\mathsf{u}^{*}+2\mathsf{u}^{*(1)})+8\bm{f}(\mathsf{u}^{*}+\mathsf{u}^{*(1)})-8\bm{f}(\mathsf{u}^{*}-\mathsf{u}^{*(1)})+\bm{f}(\mathsf{u}^{*}-2\mathsf{u}^{*(1)})\right]$
${\mathsf{s}}^{*(1)}$ $=\frac{1}{12}\left[-\text{{{s}}}(\mathsf{u}^{*}+2\mathsf{u}^{*(1)})+8\text{{{s}}}(\mathsf{u}^{*}+\mathsf{u}^{*(1)})-8\text{{{s}}}(\mathsf{u}^{*}-\mathsf{u}^{*(1)})+\text{{{s}}}(\mathsf{u}^{*}-2\mathsf{u}^{*(1)})\right]$
Acknowledgments
The work of Arpit Babbar and Praveen Chandrashekar is supported by the Department of Atomic Energy, Government of India, under project no. 12-R&D- TFR-5.01-0520.
Conflict of interest
On behalf of all authors, the corresponding author states that there is no conflict of interest.